OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Choose the correct ordered pair. y = 3x - 4 x = 9 A. (9, 23) B. (1, -1) C. (0, -4) D. (9, 31)

OpenStudy (shadowlegendx):

This is like the last one we did @Tallan

OpenStudy (e.mccormick):

Do you understand what substitution means?

OpenStudy (anonymous):

Ok so i know that 9 is x

OpenStudy (shadowlegendx):

Just input in 9 for x in the above equation

OpenStudy (shadowlegendx):

You're "substituting" in 9 for x

OpenStudy (e.mccormick):

Yep. Since x is already solved (given) it is a very simple substitution. =)

OpenStudy (shadowlegendx):

y = 3(9) - 4

OpenStudy (shadowlegendx):

Try solve for it.

OpenStudy (anonymous):

23

OpenStudy (shadowlegendx):

y = 23 is correct

OpenStudy (anonymous):

So it would be A. (9,23) right?

OpenStudy (shadowlegendx):

Yep c;

OpenStudy (shadowlegendx):

Remember to check your answers though ;p

OpenStudy (e.mccormick):

Yah, well, with one being given there is not much to check.

OpenStudy (shadowlegendx):

23 = 3(9) - 4 23 = 27 - 4 23 = 23 Correct 9 = 9 Kinda obvious ;P

OpenStudy (shadowlegendx):

@e.mccormick Just get them used to the habit of checking their work

OpenStudy (e.mccormick):

Yep. Substitution can generally be done two ways... this one is just one. If you can do it both ways and come out the same, it is (usually) good!